# Functional Stochastic Gradient MCMC for Bayesian Neural Networks

### v1

@@ 3.1 Functional SGLD for BNNs
$+(\nabla_{\mathbf{w}}f)^{T}\sigma(\mathbf{w}_{t})\mathrm{d}B_{t}.$
Given our defined functional dynamics in Equation 6 and the above relationship, the corresponding diffusion process of network weights in parameter space under our functional Langevin dynamics can be derived as
@@
$=\nabla_{\mathbf{w}}f[\nabla_{f}(\log p(\mathcal{D}|f_{t})+\log\frac{\mathrm{d}p_{0}}{\mathrm{d}r_{\mathcal{H,F}}}(f))]\mathrm{d}t+\sqrt{2}\mathrm{d}B_{t},$
@@
where the corresponding drift term $\mu(\mathbf{w}_{t})=-\nabla_{\mathbf{w}}U(f_{t})$, and the diffusion coefficient $\sigma(\mathbf{w}_{t})=\sqrt{2}\cdot\mathbf{I}_{d}$. Then, the function samples of the target posterior $p(f|\mathcal{D})$ as the stationary distribution of the functional Langevin dynamics can be obtained by discretizing and sampling the diffusion process of weights in Equation 8.
@@
In practice, the stochastic gradient variant drift term of SDE in Equation 8 is denoted as $-\nabla_{\mathbf{w}}\tilde{U}(f_{t})$, which uses the minibatch version for the likelihood term. For the gradient of the log prior distribution over functions, we firstly assume an approximated GP prior for BNNs, since in the limit of infinite width, the prior measure of BNNs converges to a GP with NNGP kernel (Neal, 1995; Lee et al., 2017; Matthews et al., 2018; Lee et al., 2019). Due to the infinite-dimensional property of functional distribution, we solve the gradient of functional prior on finite measurement points $\mathbf{X}_{\mathcal{M}}\stackrel{{\scriptstyle\text{ dif }}}{{=}}\left[\mathbf{x}_{1},\ldots,\mathbf{x}_{M}\right]^{\mathrm{T}}$ as $\nabla_{f^{\mathbf{X}_{\mathcal{M}}}}\log p(f^{\mathbf{X}_{\mathcal{M}}})$, where $f^{\mathbf{X}_{\mathcal{M}}}$ are corresponding function values evaluated at $\mathbf{X}_{\mathcal{M}}$. Then the prior $p(f^{\mathbf{X}_{\mathcal{M}}})$ is a multivariate Gaussian that could be solved analytically for the gradient. The specific discretization update rule for weight samples under functional stochastic gradient Langevin dynamics (fSGLD) is
$\mathbf{w}_{t+1}=$ $\mathbf{w}_{t}-\epsilon_{t}\nabla_{\mathbf{w}}f\Big{[}-\frac{N}{n}\sum_{i=1}^{n}\nabla_{f}\log p(y_{i}|f_{t}(x_{i};\mathbf{w}_{t}))-$ (9)
$\hskip 62.59596pt\nabla_{f^{\mathbf{X}_{\mathcal{M}}}}\log p(f_{t}^{\mathbf{X}_{\mathcal{M}}})\Big{]}+\sqrt{2\epsilon_{t}}\eta_{t}.$
@@
Similar to the naive SGLD, with the discretization step-size $\epsilon_{t}$ decays to zero, the stationary distribution for functional $p(f,t)$ of fSGLD continues to be the target posterior $p(f|\mathcal{D})$ over functions.
@@
Note that although the function samples of posterior $p(f|\mathcal{D})$ are drawn by the discretization process for corresponding diffusion of weights in Equation 9, our fSGLD is different from the existing SGLD in parameter space, which means the stationary distributions from two dynamics would be different as well. Compared to the naive parameter-space SGLD in Equation 1 controlled only by the limited parameter information, the discretization update rule for weight samples in Equation 9 under our fSGLD scheme contains information about the function mapping, which is important for accurate posterior inference. The pseudocode for fSGLD is shown in Algorithm 1 in Appendix C.
3.2 Functional SGHMC for BNNs
@@
For the naive Hamiltonian dynamics, $X=(\mathbf{w},\mathbf{z})$ is the augmentation of model parameters. We now let $X=(f,g)$ and define the Hamiltonian dynamics in function space as $H(f,g)=U(f)-\log p(g)$, where $p(g)$ is the functional auxiliary distribution induced by the distribution of auxiliary variable $\mathbf{z}$. We design the functional Hamiltonian dynamics for $f(\cdot;\mathbf{w})$ and $g(\cdot;\mathbf{z})$ as
$\mathrm{d}f_{t}(\cdot;\mathbf{w})=-(\nabla_{\mathbf{w}}f)^{T}\nabla_{\mathbf{z}}g\nabla_{g}\log p(g_{t})\mathrm{d}t,$ (10)
$\mathrm{d}g_{t}(\cdot;\mathbf{z})=-(\nabla_{\mathbf{z}}g)^{T}\nabla_{\mathbf{w}}f\nabla_{f}U(f_{t})\mathrm{d}t.$
@@
The (marginal) stationary distribution of the functional Hamiltonian dynamics defined in Equation 10 is the target posterior distribution over functions $p(f|\mathcal{D})$.
@@
The proof is given in Appendix B. ∎
Similar to the fSGLD, we also transform the above function-space Hamiltonian dynamics to its corresponding weight-space dynamics of $\mathbf{w}$ and $\mathbf{z}$ as follows:
$\mathrm{d}\mathbf{w}_{t}$ $=\frac{\partial H}{\partial\mathbf{z}}=-\nabla_{\mathbf{z}}g\cdot\nabla_{g}\log p(g_{t})\mathrm{d}t,$ (11)
$\mathrm{d}\mathbf{z}_{t}$ $=-\frac{\partial H}{\partial\mathbf{w}}=-\nabla_{\mathbf{w}}f\cdot\nabla_{f}U(f_{t})\mathrm{d}t,$
@@
$\mathbf{w}_{t+1}=$ $~{}\mathbf{w}_{t}-\epsilon_{t}\nabla_{\mathbf{z}}g\cdot\nabla_{g}\log p(g_{t})$ (12)
$\mathbf{z}_{t+1}=$ $~{}\mathbf{z}_{t}-\epsilon_{t}\nabla_{\mathbf{w}}f\cdot\nabla_{f}\tilde{U}(f_{t})$
$\hskip 14.22636pt+\epsilon_{t}C\nabla_{\mathbf{z}}g\cdot\nabla_{g}\log p(g_{t})+\sqrt{2C\epsilon_{t}}\eta_{t}.$
The stationary distribution of such functional stochastic gradient Hamiltonian is still the target functional posterior $p(f|\mathcal{D})$ (see Appendix B). Like the discussion in the last section, the diffusion process for $\mathbf{w}$ and $\mathbf{z}$ under fSGHMC scheme in Equation 12 containing abundant functional information is obviously different from the naive parameter-space SGHMC in Equation 2. The pseudocode for fSGHMC is shown in Algorithm 2 in Appendix C.
For our functional SGMCMC, the predictive distribution is obtained from the following integration process:
$p(y^{*}|x^{*},\mathcal{D})$ $=\int p(y^{*}|f(x^{*}))p(f|\mathcal{D})\mathrm{d}f$ (13)
@@
where function draws of posterior over functions are generated from corresponding samples of weight $\mathbf{w}^{(j)}(j=1,2,...,S)$ using fSGLD or fSGHMC.
@@ 4 Related Work
Functional Variational Inference for BNNs. With unresolved concerns about parameter-space weight prior to posterior approximation in BNNs, Bayesian inference in function space has attracted increasing attention. In order to specify meaningful functional prior like Gaussian processes, Sun et al. (2019) proposed a kind of functional variational Bayesian neural networks by maximizing the functional evidence lower bound (ELBO) to optimize the Kullback–Leibler (KL) divergence between the posterior processes and variational posterior over functions explicitly. They solved the KL divergence between infinite-dimensional stochastic processes on a finite marginal measurement set and used the Spectral Stein gradient estimator for implicit functional prior and variational posterior over functions of BNNs (Shi et al., 2018). However, Burt et al. (2020) pointed out some limitations of employing KL divergence in infinite-dimensional function space. The variational posterior should be absolutely continuous with respect to the prior measure to guarantee the existence of the Radon-Nikodym derivatives involved in the definition of functional KL divergence (Matthews et al., 2016; Wild et al., 2022). Otherwise, for example, the KL divergence between distributions over functions generated from two BNNs with different network architectures is ill-defined and could be infinite (Ma & Hernández-Lobato, 2021). To avoid this issue, Rudner et al. (2022) proposed to approximate the distributions over functions of a BNN prior and variational posterior that share the same network structures as two Gaussian distributions via the linearization of their mean parameters and derived a tractable and well-defined functional variational objective for BNNs. Considering the potential weaknesses of KL divergence for stochastic processes, Tran et al. (2022) proposed to match a BNN prior to a GP prior by minimizing the 1-Wasserstein distance (Kantorovich, 1960) to obtain a more interpretable functional prior for BNNs.
@@
Markov Chain Monte Carlo for BNNs. Papamarkou et al. (2022) summarized the challenges in applying MCMC approaches to posterior inference for BNNs in four aspects: computational cost, model structure, weight symmetries, and prior specification. MCMC methods have not been broadly adopted in posterior inference for large modern BNNs until the introduction of scalable Monte Carlo algorithms, such as the stochastic gradient MCMC (Welling & Teh, 2011; Chen et al., 2014; Ahn et al., 2014; Garriga-Alonso & Fortuin, 2021). Ma et al. (2015) designed a unifying framework for stochastic gradient MCMC whereby all past dynamics-based samplers and their minibatch variants can be cast into it. For the multimodal properties of posterior, Zhang et al. (2020) proposed a cyclical step size schedule in stochastic gradient MCMC for BNNs to improve the sampling efficiency by which they proved the non-asymptotic convergence of their method. On the other hand, Cobb & Jalaian (2021) emphasized the limitations of the use of stochastic gradients in HMC (Shahbaba et al., 2014), and proposed a scaling full HMC for BNNs by a symmetric splitting integration scheme. And Izmailov et al. (2021) also used full-batch HMC to investigate the foundational properties of posterior for BNNs. Alexos et al. (2022) proposed a structured stochastic gradient MCMC that combines structure variational inference and MCMC to improve the slow mixing in high dimensional BNNs.
@@
Figure 1: 1-D extrapolation example. The green line is the ground true function, and the blue lines correspond to the mean of samples from posterior predictions. Black dots denote 20 training points; shadow areas represent the predictive standard deviations.
5 Experiments
We evaluate the predictive performance and uncertainty quantification of our functional stochastic gradient MCMC methods on several benchmark tasks, including a synthetic extrapolation example, multivariate regressions on UCI datasets, image classification tasks, and contextual bandits. We compare fSGLD and fSGHMC with naive SGLD, SGHMC, and competing functional variational inference methods to illustrate the sampling efficiency of our functional stochastic gradient MCMC methods.
5.1 Extrapolation on Synthetic Data
To evaluate the fitting ability and uncertainty estimation of our method, we first consider a 1-D oscillation curve extrapolation on a synthetic dataset. For input, we randomly sampled half of 20 observation points from Uniform$(-0.75,-0.25)$, and the other half are from Uniform$(0.25,0.75)$. The output is modeled through the polynomial function: $y=\sin(3\pi x)+0.3\cos(9\pi x)+0.5\sin(7\pi x)+\epsilon$ with noise $\epsilon\sim\mathcal{N}(0,0.5^{2})$. We compared our fSGLD and f SGHMC with naive SGLD, SGHMC, and two representative functional variational methods: FBNN (Sun et al., 2019) and IFBNN (Wu et al., 2023). The model is a fully-connected neural network with two hidden layers. For our functional fSGLD, fSGHMC, and the two functional variational inference methods, we use the same functional GP prior with the RBF kernel pre-trained on the input dataset. There are 40 measurement points for the approximate estimation of the gradient of log functional prior in fSGLD and fSGHMC, which are randomly sampled from training data and an additional 40 inducing points from Uniform $(-1,1)$. For all sampling methods, we use 2000 burn-in iterations and 8000 iterations for 80 samples (to reduce correlations between samples, we draw separated samples every 100 epochs). Functional variational inference methods are trained for 10000 epochs for fair comparison. Results are shown in Figure 1. Figure 1(a) and Figure 1(b) show that fSGLD and fSGHMC can recover the key characteristic of the curve in the observation range and provide reasonable uncertainty estimations in the unseen region. fSGHMC performs slightly better than fSGLD. In contrast, for naive SGMCMC methods, Figure 1(d) illustrates that SGLD underestimates the uncertainty in non-observation areas, especially evident in the leftmost non-observed $[-1,-0.75]$ interval. Meanwhile, Figure 1(e) shows that SGHMC suffers from some unreasonable manifestations of the uncertainty estimation, such as the undeserved uncertainty expansion in the interval $[-0.4,-0.25]$ of the input data range. On the other hand, the rightmost column shows results from two functional variational methods. We found that FBNN in Figure 1(c) is unable to fit the target curve in the training data range $[-0.75,0.25]$ and performs poorly for uncertainty quantification. Figure 1(f) shows that IFBNN also severely underestimates the uncertainty, which is consistent with the findings of Ormerod & Wand (2010) and Zhang et al. (2018) that variational approximation typically underestimates the posterior variance.

### v2

@@ 3.1 Functional Langevin Dynamics for BNNs
$+(\nabla_{\mathbf{w}}f)^{T}\sigma(\mathbf{w}_{t})\mathrm{d}B_{t},$
where $\mathrm{d}\mathbf{w}_{t}=\mu(\mathbf{w}_{t})\mathrm{d}t+\sigma(\mathbf{w}_{t})\mathrm{d}B_{t}$ is the general diffusion form for $\mathbf{w}$. By combining our functional Langevin dynamics defined in Equation 3 with this differential relationship between $f(\cdot;\mathbf{w})$ and $\mathbf{w}$, it is straightforward to reversely derive the corresponding diffusion process for $\mathbf{w}$ in parameter spaces as follows:
@@
$=-\nabla_{\mathbf{w}}f_{t}\left[\nabla_{f}(-\log p(\mathbf{Y}_{\mathcal{D}}|f_{t}(\mathbf{X}_{\mathcal{D}};\mathbf{w}))+I_{0}(f_{t}))\right]\mathrm{d}t$
$+\sqrt{2}\mathrm{d}B_{t},$
@@
where the drift term $\mu(\mathbf{w})=-\nabla_{\mathbf{w}}U(f)$ , and the diffusion coefficient $\sigma(\mathbf{w}_{t})=\sqrt{2}\cdot\mathbf{I}_{k}$. Then, by discretizing and sampling from this corresponding diffusion process for the network parameters, we can obtain function samples from the target posterior measure $P_{f|\mathcal{D}}$ in a tractable manner.
@@
In practice, the stochastic gradient variant for the drift term of SDE in Equation 5 is denoted as $-\nabla_{\mathbf{w}}\tilde{U}(f)$, which utilizes the minibatch computation for the likelihood term. To estimate the gradient of the prior measure, we assume an approximate GP prior for BNNs. This is motivated by the fact that, in the limit of infinite width, the prior measure of BNNs converges to a GP with the Neural Network Gaussian Process (NNGP) kernel (Neal,, 1995; Lee et al.,, 2017; Matthews et al.,, 2018; Lee et al.,, 2019). Due to the infinite-dimensional nature of the functional distribution, we solve the gradient of functional prior on finite measurement points $\mathbf{X}_{\mathcal{M}}\stackrel{{\scriptstyle\text{ def }}}{{=}}\left[\mathbf{x}_{1},\ldots,\mathbf{x}_{M}\right]^{\mathrm{T}}$ as $\nabla_{f^{\mathbf{X}_{\mathcal{M}}}}\log P_{0}(f^{\mathbf{X}_{\mathcal{M}}})$, where $f^{\mathbf{X}_{\mathcal{M}}}$ are corresponding function values evaluated at $\mathbf{X}_{\mathcal{M}}$. The prior $P_{0}(f^{\mathbf{X}_{\mathcal{M}}})$ reduces to a multivariate Gaussian, allowing for an analytical solution for the gradient. The specific discretization update rule for network parameters samples under functional stochastic gradient Langevin dynamics (fSGLD) is given by
$\mathbf{w}_{t+1}=$ $\mathbf{w}_{t}-\epsilon_{t}\nabla_{\mathbf{w}}f_{t}\Big{[}-\frac{N}{n}\sum_{i=1}^{n}\nabla_{f}\log p(y_{i}|f_{t}(x_{i};\mathbf{w}_{t}))-$ (6)
$\hskip 62.59596pt\nabla_{f^{\mathbf{X}_{\mathcal{M}}}}\log P_{0}(f_{t}^{\mathbf{X}_{\mathcal{M}}})\Big{]}+\sqrt{2\epsilon_{t}}\eta_{t}.$
@@
Like the standard SGLD, with the discretization step-size $\epsilon_{t}$ decays to zero, the stationary measure of fSGLD continues to be the target posterior $P_{f|\mathcal{D}}$ over functions.
@@
Note that in contrast to the naive parameter-space SGLD in Equation 1, which is driven solely by limited parameter information, the discretization update rule for samples of network parameters in Equation 6 under our fSGLD scheme incorporates information about the function mapping. This additional functional information plays a crucial role in achieving more accurate posterior inference. The pseudocode for fSGLD is shown in Algorithm 1 in Appendix D.
3.2 Functional Hamiltonian dynamics for BNNs
@@
For the naive Hamiltonian dynamics, $X=(\mathbf{w},\mathbf{z})$ represents the augmentation of model parameters. We now let $X=(f,g)$ and define the Hamiltonian dynamics in function space as $H(f,g)=U(f)-\log p(g)$, where $p(g)$ is the functional auxiliary probability measure induced by the distribution of auxiliary variable $\mathbf{z}$. We then design the functional Hamiltonian dynamics for $f(\cdot;\mathbf{w})$ and $g(\cdot;\mathbf{z})$ as
$\mathrm{d}f_{t}(\cdot;\mathbf{w})=-(\nabla_{\mathbf{w}}f_{t})^{T}\nabla_{\mathbf{z}}g_{t}\nabla_{g}\log p(g_{t})\mathrm{d}t,$ (7)
$\mathrm{d}g_{t}(\cdot;\mathbf{z})=-(\nabla_{\mathbf{z}}g_{t})^{T}\nabla_{\mathbf{w}}f_{t}\nabla_{f}U(f_{t})\mathrm{d}t.$
@@
The (marginal) stationary probability measure of the functional Hamiltonian dynamics defined in Equation 7 is the target posterior over functions $P_{f|\mathcal{D}}$.
@@
The proof is given in Appendix C. ∎
Similar to the fSGLD, we also transform the above function-space Hamiltonian dynamics to its corresponding parameter-space dynamics of $\mathbf{w}$ and $\mathbf{z}$ as follows:
$\mathrm{d}\mathbf{w}_{t}$ $=\frac{\partial H}{\partial\mathbf{z}}=-\nabla_{\mathbf{z}}g_{t}\cdot\nabla_{g}\log p(g_{t})\mathrm{d}t,$ (8)
$\mathrm{d}\mathbf{z}_{t}$ $=-\frac{\partial H}{\partial\mathbf{w}}=-\nabla_{\mathbf{w}}f_{t}\cdot\nabla_{f}U(f_{t})\mathrm{d}t,$
@@
$\mathbf{w}_{t+1}=$ $~{}\mathbf{w}_{t}-\epsilon_{t}\nabla_{\mathbf{z}}g_{t}\cdot\nabla_{g}\log p(g_{t})$ (9)
$\mathbf{z}_{t+1}=$ $~{}\mathbf{z}_{t}-\epsilon_{t}\nabla_{\mathbf{w}}f_{t}\cdot\nabla_{f}\tilde{U}(f_{t})$
$\hskip 14.22636pt+\epsilon_{t}C\nabla_{\mathbf{z}}g_{t}\cdot\nabla_{g}\log p(g_{t})+\sqrt{2C\epsilon_{t}}\eta_{t}.$
The stationary probability measure of such functional stochastic gradient Hamiltonian can still be guaranteed as the target functional posterior $P_{f|\mathcal{D}}$ (see Appendix C). Like the discussion in the last section, the diffusion process for $\mathbf{w}$ and $\mathbf{z}$ under fSGHMC scheme in Equation 9 incorporates abundant functional information is obviously different from the naive parameter-space SGHMC in Equation 2. The pseudocode for fSGHMC is shown in Algorithm 2 in Appendix D.
For our functional MCMC schemes, the predictive distribution is obtained from the following integration process:
$p(y^{*}|x^{*},\mathcal{D})$ $=\int p(y^{*}|f(x^{*}))P_{f|\mathcal{D}}\mathrm{d}f$ (10)
@@
where function draws of posterior over functions are generated from corresponding samples of network parameters $\mathbf{w}^{(j)}(j=1,2,...,S)$ using fSGLD or fSGHMC.
@@ 4 Related Work
Functional Variational Inference for BNNs. To specify meaningful functional priors like Gaussian processes, Sun et al., (2019) proposed a functional evidence lower bound (ELBO) to optimize the Kullback–Leibler (KL) divergence between variational posterior and the true posterior in function space explicitly. They solved the KL divergence between infinite-dimensional stochastic processes on finite marginal measurement sets. Due to the analytical intractability of the supremum over marginal KL divergences in functional ELBO, Rudner et al., (2022) proposed to approximate the posterior and prior over functions of BNNs as Gaussian distributions through local linearization and derived a more tractable functional variational objective for BNNs. Considering the potential weaknesses of KL divergence for stochastic processes Burt et al., (2020), Tran et al., (2022) proposed to match a BNN prior to a GP prior by minimizing the 1-Wasserstein distance (Kantorovich,, 1960) to obtain a more interpretable functional prior.
@@
Markov Chain Monte Carlo for BNNs. Papamarkou et al., (2022) summarized the challenges in applying MCMC approaches to posterior inference for BNNs in four aspects: computational cost, model structure, weight symmetries, and prior specification. MCMC methods had limited success in being broadly adopted for posterior inference in large modern BNNs until the introduction of scalable Monte Carlo algorithms, such as the stochastic gradient MCMC (Welling and Teh,, 2011; Chen et al.,, 2014; Ahn et al.,, 2014; Garriga-Alonso and Fortuin,, 2021). Ma et al., (2015) designed a unifying framework that casts all existing dynamics-based samplers and their minibatch variants into it. Regarding the multimodal properties of posterior distributions, Zhang et al., (2020) proposed a cyclical step-size schedule in stochastic gradient MCMC for BNNs to improve the sampling efficiency. On the other hand, Cobb and Jalaian, (2021) emphasized the limitations of stochastic gradients in HMC and proposed a scaling full-batch HMC for BNNs using a symmetric splitting integration scheme. However, all these MCMC methods are performed in parameter space.
5 Experiments
We evaluate the predictive performance and uncertainty quantification of our functional stochastic gradient MCMC schemes on several benchmark tasks, including a synthetic extrapolation example, multivariate regressions on UCI datasets, image classification tasks, and contextual bandits. We compare fSGLD and fSGHMC with naive SGLD, SGHMC, and competing functional variational inference methods to illustrate the sampling efficiency of our functional stochastic gradient MCMC schemes.
5.1 Extrapolation on Synthetic Data
To evaluate the fitting ability and uncertainty estimation, we first consider a 1-D oscillation curve extrapolation on a synthetic dataset. For input, we randomly sampled half of 20 observation points from Uniform$(-0.75,-0.25)$, and the other half are from Uniform$(0.25,0.75)$. The output is modeled through the polynomial function: $y=\sin(3\pi x)+0.3\cos(9\pi x)+0.5\sin(7\pi x)+\epsilon$ with noise $\epsilon\sim\mathcal{N}(0,0.5^{2})$. We compared our fSGLD and fSGHMC with naive SGLD, SGHMC, and two representative functional variational methods: FBNN (Sun et al.,, 2019) and IFBNN (Wu et al.,, 2023). The model is a fully-connected neural network with two hidden layers. For our functional fSGLD, fSGHMC, and the two functional variational inference methods, we use the same functional GP prior with the RBF kernel pre-trained on the input dataset. There are 40 measurement points for the approximate estimation of the gradient of log functional prior in fSGLD and fSGHMC, which are randomly sampled from training data and an additional 40 inducing points from Uniform $(-1,1)$. For all sampling methods, we use 2000 burn-in iterations and 8000 iterations for 80 samples (to reduce correlations between samples, we draw separated samples every 100 epochs). Functional variational inference methods are trained for 10000 epochs for fair comparison. Results are shown in Figure 1. Figure 1(a) and Figure 1(b) show that fSGLD and fSGHMC can recover the key characteristic of the curve in the observation range and provide reasonable uncertainty estimations in the unseen region. fSGHMC performs slightly better than fSGLD. In contrast, for naive SGMCMC methods, Figure 1(d) illustrates that SGLD underestimates the uncertainty in non-observation areas, especially evident in the leftmost non-observed $[-1,-0.75]$ interval. Meanwhile, Figure 1(e) shows that SGHMC suffers from some unreasonable manifestations of the uncertainty estimation, such as the undeserved uncertainty expansion in the interval $[-0.4,-0.25]$ of the input data range. On the other hand, the rightmost column shows results from two functional variational methods. We found that FBNN in Figure 1(c) is unable to fit the target curve in the training data range $[-0.75,0.25]$ and performs poorly for uncertainty quantification. Figure 1(f) shows that IFBNN also severely underestimates the predictive uncertainty, which is consistent with the findings of Ormerod and Wand, (2010) and Zhang et al., (2018) that variational approximation typically underestimates the posterior variance. More analysis of mixing rate and computational complexity are shown in Appendix F and Appendix G, respectively. Appendix H presents a wide ablation study about the sample size, measurement points and functional prior.
@@ 5.1 Extrapolation on Synthetic Data
Figure 1: 1-D extrapolation example. The green line is the ground true function, and the blue lines correspond to the mean of samples from posterior predictions. Black dots denote 20 training points; shadow areas represent the predictive standard deviations. For more details, see Appendix E.